OpenStudy (anonymous):

If h(2) = 3 and h'(2) = −6, find

myininaya (myininaya):

tangent line at x=2?

OpenStudy (anonymous):

\[\frac{ d }{ dx }\(\frac{ h(x) }{ x }) x=2\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

are you sure? what is that you are writing?

zepdrix (zepdrix):

Did you mean to write h(x) / x ? Looks like too many x's 0_o

myininaya (myininaya):

\[\frac{d}{dx} (\frac{h(x)}{x})=2\]?

OpenStudy (anonymous):

ill take a screen shot

myininaya (myininaya):

do you know how to use the quotient rule or even the product rule will work (if we rewrite it as a product)

OpenStudy (anonymous):

i have it in my notes but i don't feel confident enough to do it on my own

myininaya (myininaya):

\[\frac{d}{dx}(h(x)x^{-1})=(h(x))'(x^{-1})+(h(x)) (x^{-1})' \text{ by product rule } \]

myininaya (myininaya):

find the derivative of x^(-1) derivative of h is easy it is h'

OpenStudy (anonymous):

what would our first step be then for solving the problem

myininaya (myininaya):

finding the derivative of x^(-1)

OpenStudy (anonymous):

okay how do we do that then.. lol

myininaya (myininaya):

power rule

OpenStudy (anonymous):

so then move the number over pass the x and then subtract exponents by 1

myininaya (myininaya):

yes and that gives you?

myininaya (myininaya):

\[\frac{d}{dx}(h(x)x^{-1})=(h(x))'(x^{-1})+(h(x)) (x^{-1})' \text{ by product rule } \] when you are done you can replace the (x^(-1))' with what you fund for it

myininaya (myininaya):

found*

myininaya (myininaya):

once you found the derivative your next is to replace the x's with 2's and to recall what h(2) and h'(2) was

myininaya (myininaya):

and then order of operations